1. Jamille walks from Julie Bell's house 20 blocks west and 10 blocks north to the house of Meredam. She then walks 11 blocks west and 2 blocks north to Jm's house. How far away is Julie Bell's house from Jm's house?


2. There are 15 scholarship to be apportioned among 238 English majors,502 Psychology majors,and 355 Math majors. Determine the standard divisor


1
Expert's answer
2022-01-11T19:17:13-0500

1. We have two perpendicular directions.

In direction west-east Jamille walks  20 blocks west and 11 blocks west


"20+11=31"

In direction south-north Jamille walks  10 blocks north and 2 blocks north


"10+2=12"

In total Jamille walks from Julie Bell's house 31 blocks west and 12 blocks north to the Jm's house.

Therefore Julie Bell's house is 12 south and 31 blocks east from Jm's house.


2. Since there is a total population of "238+502+355=1095" and "15" seats. The standard divisor will be 


"\dfrac{1095}{15}=73"